ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

12. Probability and Statistics

1991 MS Classification: 62, 92 Postelnicu, Tiberiu, Centre for Mathematical Statistics, Bucharest, Romania Specific Statistical Method for Binary Data A new clustering method for binary data is proposed. The method is based on the concept of “homogeneity” within a set of two or more Operational Taxonomic Units, The paper proposes and studies homogeneities for binary and for ordered multi-states data. The “h” and “H” homogeneities are presented and an application in biology is given. The homogeneity denoted by “h” has the remarkable quality of generalizing the famous JACCARD’s similarity coefficient. Moreover, a probability value is associated with every possible cluster and only statistically significant are considered. Reference: J. W. Drane, T. Postelnicu and C. Y. Fu, Theorems on binary responses to combination of stimuli, Biom. Journal,34 (1992), 605 - 615.